A solution is prepared by dissolving 20g benzene in 100g carbon tetrachloride, what is the % by mass of the solution?

a. 17%
b. 20%
c. 83%
d. 500%
1
Expert's answer
2020-05-01T14:31:45-0400

Solution:

m(C6H6) = 20 g

m(CCl4) = 100 g

m(solution) = m(C6H6) + m(CCl4) = 20g + 100g = 120 g

w(C6H6) = m(C6H6)/m(solution)*100% = 20g/120g*100% = 17%

Answer: 17%